### v2

# Infer-and-widen, or not?

Abstract
In recent years, there has been substantial interest in the task of selective inference: inference on a parameter that is selected from the data. Many of the existing proposals fall into what we refer to as the infer-and-widen framework: they produce symmetric confidence intervals whose midpoints do not account for selection and therefore are biased; thus, the intervals must be wide enough to account for this bias. In this paper, we investigate infer-and-widen approaches in three vignettes: the winner’s curse, maximal contrasts, and inference after the lasso. In each of these examples, we show that a state-of-the-art infer-and-widen proposal leads to confidence intervals that are much wider than simple alternatives when all methods are tuned to yield identical randomized selection events. Furthermore, even an “oracle” infer-and-widen confidence interval — the narrowest possible interval that could be theoretically attained via infer-and-widen — is often wider than these alternatives.
1 Introduction
In classical statistics, the target of inference — for instance, the null hypothesis that we wish to test, or the parameter for which we wish to construct a confidence interval — is fixed. However, motivated by the explosion of data across a variety of fields, recent interest has focused on the setting where the target of inference is itself a function of the data, and therefore is random. The terms selective inference and post-selection inference describe the set of strategies available for conducting inference on a random target.
To formalize this, consider a random variable $Y$ with domain $\mathcal{Y}$, and a collection of parameters $\theta$ whose elements can be indexed by a (possibly infinite) index set $\mathcal{I}$. For a fixed index $i\in\mathcal{I}$, and for $\alpha\in(0,1)$, a classical confidence interval for $\theta_{i}$ is a confidence interval $\mathrm{CI}_{i}^{\alpha}(Y)$ with $1-\alpha$ coverage in the sense that
$\mathbb{P}\left(\theta_{i}\not\in\mathrm{CI}_{i}^{\alpha}(Y)\right)\leq\alpha.$ (1.1)
Now, suppose that $i\in\mathcal{I}$ is no longer a fixed index; instead, some selection rule ${\mathcal{S}}:\mathcal{Y}\to\mathcal{I}$ uses the data to select a parameter on which to conduct inference. A confidence interval provides valid unconditional selective coverage (Fithian et al., 2017) at level $\alpha$ if
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{\alpha}(Y)\right)\leq\alpha.$ (1.2)
Unconditional coverage is not guaranteed by classical confidence intervals. In this paper, we consider methods that yield unconditional guarantees, in the sense of 1.2. Some of these confidence intervals also yield (stronger) conditional selective guarantees, in the sense that
$\mathbb{P}\left(\theta_{i}\not\in\mathrm{CI}_{i}^{\alpha}(Y)~{}|~{}\mathcal{C}(Y)\right)\leq\alpha$ (1.3)
for an event $\mathcal{C}(Y)$ satisfying $\sigma\left(\mathcal{C}(Y)\right)\supseteq\sigma\left({\mathcal{S}}(Y)\right)$, where $\sigma(\cdot)$ denotes the sigma algebra generated by a random variable, i.e., the coverage guarantee in 1.3 conditions on at least the selection event; by the law of total expectation, conditional coverage 1.3 implies unconditional coverage 1.2.
1.1 The infer-and-widen framework
As presented in Box 1.1, the infer-and-widen framework involves (Step 1) defining a target of inference using some or all of the data, and then (Step 2) symmetrically adjusting the critical values arising from “classical” inference to account for the fact that the target of inference is not fixed in order to achieve unconditional coverage (1.2). The simplest examples of this strategy are the Bonferroni (Bonferroni, 1936), Holm-Bonferroni (Holm, 1979), and Šidák (Šidák, 1967) procedures, which are suitable when a finite number of targets of inference are considered (and in the case of Šidák, under additional assumptions). In the special case of inference on contrasts of a linear model, the Scheffé correction is available (Scheffé, 1959); this idea is further refined by the PoSI method (Berk et al., 2013). While conceptually attractive, these approaches tend to be quite conservative (Bonferroni, Holm-Bonferroni, Šidák), are limited to a linear model setting (Scheffé, PoSI), or are computationally intractable (PoSI).
Other recent infer-and-widen approaches have attempted to address these shortcomings. The local simultaneous inference approach of Zrnic and Fithian (2024b) improves upon the Bonferroni correction by focusing on the subset of hypotheses that are most likely given the observed data. By contrast, the algorithmic stability proposal of Zrnic and Jordan (2023) adapts to the details of the selection algorithm. By drawing connections with the field of differential privacy (Dwork and Roth, 2014), the “stability” of the approach for selecting the target of inference in Step 1 of Box 1.1 is quantified; this, in turn, leads to a stability-specific adjustment of the critical values in Step 2 of Box 1.1. In a sense, algorithmic stability achieves the lofty goals laid out in the PoSI proposal of Berk et al. (2013) by creating a framework to compute a selection-specific adjustment to the critical values (see also the online supplement).
[!htb] Box 1: Infer-and-widen. 1. A target of inference is selected from the data. 2. “Classical” inference is conducted on all of the data, with a symmetric adjustment of the critical values to account for the data-driven selection of the target of inference.
1.2 Other approaches for valid selective inference
Not all approaches that yield unconditional guarantees of the form (1.2) fall into the infer-and-widen framework. Some methods produce asymmetric confidence intervals around the biased point estimate: for instance, Benjamini et al. (2019) and Fuentes et al. (2018) construct an asymmetric Bonferroni correction that accounts for the direction of the bias. More recently, Zrnic and Fithian (2024a) extend the local simultaneous approach of Zrnic and Fithian (2024b) by additionally adapting the correction to the specific selection algorithm. The hybrid approaches of Andrews et al. (2019) and McCloskey (2024) apportion a fraction of the Type 1 error budget to rule out unlikely selection events, and then construct a conditionally-valid confidence interval on the likely hypotheses; note that this only guarantees unconditional coverage 1.2.
Furthermore, a number of methods yield coverage guarantees conditional on the selection event, i.e., (1.3). Some methods do this by obtaining independent training and test sets; then, the training set can be used to define the target of inference, and the test set to conduct inference on this target. Because the two sets are independent, we can conduct inference using the test set as though the target of inference were fixed, despite the fact that it is actually a function of the training data. Sample splitting, applicable in the setting of independent and identically distributed observations, provides one approach to obtain independent training and test sets (Cox, 1975). More recent data thinning proposals generalize the idea of sample splitting, and can be applied when only a single observation is available (Rasines and Young, 2023; Neufeld et al., 2024; Dharamshi et al., 2024).
In cases where it is either impossible or undesirable to create independent training and test sets, strategies that achieve conditional guarantees of the form 1.3 may still be available. For instance, a target of inference can be defined using all of the data, and then inference can be conducted conditional on the event that the data led to this particular target of inference. Examples include the proposals of Fithian et al. (2017); Lee et al. (2016); Chen and Witten (2023); Gao et al. (2024); Andrews et al. (2019). Alternatively, if a researcher is willing to define the target of inference using only a subset of the data or using a noisy version of the data, then data fission Leiner et al. (2023) or randomized conditional selective inference Tian and Taylor (2018); Panigrahi et al. (2024) may be applicable.
1.3 Contributions and organization
In this paper, we conduct an empirical and theoretical evaluation of infer-and-widen strategies. Since randomization is known to improve the performance of selective inference methods (Panigrahi et al., 2024; Tian and Taylor, 2018), we only consider randomized infer-and-widen approaches. The bottom line of our paper is as follows: when comparing infer-and-widen strategies to alternative selective inference methods that make use of identical selection events and yield conditional 1.3 (rather than unconditional 1.2 guarantees), the latter often outperform the former in terms of confidence interval width. Furthermore, the latter often outperform even the “oracle” infer-and-widen procedure, the narrowest possible interval that could be theoretically attained within the infer-and-widen framework.
The rest of this paper is organized as follows. In Section 2, we formalize the infer-and-widen framework. In Sections 3, 4, and 5, we compare its performance to alternative approaches in two vignettes considered in Zrnic and Jordan (2023), and in an example involving model selection via the lasso. In Section 6, we show analytically that the infer-and-widen interval can suffer from a highly biased midpoint. We close with a discussion in Section 7. Theoretical results are proven in the online supplement. Scripts to fully reproduce all numerical results can be found at https://github.com/rflperry/infer_and_widen.
2 Formalizing the infer-and-widen framework
We now formalize the infer-and-widen (I&W) framework, outlined in Box 1.1. This involves a modification to the classical confidence interval $\mathrm{CI}_{i}^{\alpha}$ in (1.1) to account for the fact that $\theta_{\mathcal{S}(Y)}$ is a data-driven parameter. Based on the particulars of the selection function $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ and data $Y$, a modified $\alpha$-level, $H(\alpha,{\mathcal{S}},Y)$, is derived so that $\mathrm{CI}^{H(\alpha,{\mathcal{S}},Y)}_{\mathcal{S}(Y)}(Y)$ is unconditionally valid at level $\alpha$ in the sense of 1.2, i.e.,
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{H(\alpha,{\mathcal{S}},Y)}(Y)\right)\leq\alpha.$ (2.1)
In effect, this procedure symmetrically widens the classical confidence interval to account for bias in its midpoint.
In a Bonferroni correction we have $H(\alpha,{\mathcal{S}},Y)=\alpha/|\mathcal{I}|$, and in the Holm-Bonferroni correction it is slightly larger. Since these approaches adapt neither to the selection rule $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ nor to the data $Y$, the constant $H(\alpha,{\mathcal{S}},Y)$ is easy to compute but overly conservative.
The foundational PoSI proposal (Berk et al., 2013) considers the setting of linear regression with fixed regressors. They derive a constant $H(\alpha,{\mathcal{S}},Y)$ that enables inference on the coefficient(s) in any submodel, regardless of the form of $\mathcal{S}(\cdot)$ or observed data $Y$. However, for a particular selection rule $\mathcal{S}(\cdot)$, the resulting confidence interval may be quite conservative, and computing the constant $H(\alpha,{\mathcal{S}},Y)$ is challenging or infeasible.
Despite the notation $H(\alpha,{\mathcal{S}},Y)$ used in (2.1), the aforementioned infer-and-widen proposals result in confidence intervals whose widths do not depend on the particulars of the selection rule $\mathcal{S}(\cdot)$ nor data $Y$: that is, $H(\alpha,{\mathcal{S}},Y)$ could be written more simply as $H(\alpha)$.
In contrast are two newer approaches. The local simultaneous inference method of Zrnic and Fithian (2024b) apportions the Type-1 error budget $\alpha$, first to use the data $Y$ to rule out the most unlikely hypotheses, and then to perform a Bonferroni correction over a subset of all $\left|\mathcal{I}\right|$ hypotheses. We note that in this case, $H(\alpha,{\mathcal{S}},Y)$ could be written more simply as $H(\alpha,Y)$, as the correction does not take into account the selection rule. On the other hand, the algorithmic stability proposal of Zrnic and Jordan (2023) provides constants $H(\alpha,{\mathcal{S}})$ that are tailored to the particular selection rule ${\mathcal{S}}(\cdot)$ under consideration, although not to the data $Y$. The online supplement contains an overview of the algorithmic stability proposal; intuitively, a more “stable” selection rule ${\mathcal{S}}(\cdot)$ requires a less substantial adjustment to the critical values in Step 2 of Box 1.1.
We close with a remark about the infer-and-widen framework: at first glance, it appears that an “advantage” of infer-and-widen, over alternative selective inference approaches, is that the interval in (2.1) is of the same form as the classical interval in the absence of selection (1.1), i.e., no bespoke methods for computing a confidence interval are required. However, this is a double-edged sword, in the sense that bias induced by the selection event is perpetuated into the confidence interval. Thus, an infer-and-widen confidence interval that attains the nominal coverage may be extremely wide, to account for bias in the interval’s midpoint. We will see that this is a major drawback of infer-and-widen.
3 Vignette #1: The winner’s curse
Suppose that we observe $n$ independent Gaussian random variables $Y_{1},\ldots,Y_{n}$ with common known variance, and are interested in conducting inference on the unknown mean of the largest observed variable. That is,
$Y\sim{\bf{\rm N}}\,_{n}(\mu,\sigma^{2}I_{n}),$ (3.1)
the selection rule is
$\mathcal{S}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}Y_{i},$ (3.2)
and our target of inference is $\mu_{\mathcal{S}(Y)}$. It is not hard to see that a $1-\alpha$ confidence interval for $\mu_{\mathcal{S}(Y)}$ that treats $\mathcal{S}(Y)$ as though it were fixed (rather than a function of the data) will fail to achieve the nominal coverage. This is the “winner’s curse” (Andrews et al., 2019).
Figure 1: Vignette #1 under (3.1) with $\mu=0$, $\sigma=1$, and the selection rule in 3.3. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of the infer-and-widen 3.4 and data fission (DF) 3.7 intervals with $95\%$ coverage across values of the sample size $n$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. (b, c): Fixing $n=100$, we display the average width of the “oracle” infer-and-widen interval and the data fission interval 3.7 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval based on algorithmic stability 3.4 are also displayed. The Laplacian noise variance is $0.33$ and $3$ in panels (b) and (c), respectively.
To achieve a “stable” algorithm, Zrnic and Jordan (2023) consider a randomized version of the selection rule (3.2) via the addition of centered Laplace noise,
$\mathcal{S}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}(Y_{i}+\zeta_{i}),$ (3.3)
where $\zeta_{1},\ldots,\zeta_{n}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$ and $c$ is a positive constant that governs the amount of additive noise. When $c=0$, $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ simplifies to $\mathcal{S}(Y)$ in (3.2).
Proposition 1 (An infer-and-widen interval based on (3.3), given by (Zrnic and Jordan, 2023)).
Under 3.1, consider the selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(\cdot)$ defined in (3.3). For any $\eta>0$ and $\nu\in(0,\alpha)$ such that $\eta c\geq 2z_{1-\alpha(\alpha-\nu)/2n}$, where $z_{q}$ denotes the $q$th quantile of the standard normal distribution, the infer-and-widen confidence interval
$\mathrm{CI}_{{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\alpha+\nu)e^{-\eta}}(Y):=\left[Y_{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}\pm\sigma z_{1-\alpha(1-\alpha+\nu)e^{-\eta}/2}\right]$ (3.4)
has $1-\alpha$ unconditional selective coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 2.1.
We will now construct an alternative interval that makes use of exactly the same selection event, (3.3), using ideas from the data fission proposal of Leiner et al. (2023). Towards this end, note that $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)=\mathcal{S}(Y^{\mathrm{train}})$ where $Y^{\mathrm{train}}:=Y+\zeta$. Since $Y\mid Y^{\mathrm{train}}$ does not admit a tractable distribution, we will additionally condition on $\left({\rm sign}(\zeta_{1}),\ldots,{\rm sign}(\zeta_{n})\right)$. This yields a multivariate truncated normal distribution with independent marginals.
Proposition 2 (A data fission interval based on (3.3)).
Under (3.1), consider the selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ defined in (3.3). Define $\delta_{i}:={\rm sign}(\zeta_{i})$ and $A_{i}:=\{y_{i}\in\mathbb{R}:\delta_{i}y_{i}\leq\delta_{i}(Y_{i}+\zeta_{i})\}$. Let $\Phi_{\mu,\sigma^{2},A_{i}}(\cdot)$ denote the CDF of a ${\bf{\rm N}}\,(\mu,\sigma^{2})$ distribution truncated to the set $A_{i}$, and define $a_{i}$ and $b_{i}$ to satisfy
$\Phi_{a_{i}+\delta_{i}\sigma^{2}/c,\sigma^{2},A_{i}}(Y)=1-\alpha/2$ (3.5)
and
$\Phi_{b_{i}+\delta_{i}\sigma^{2}/c,\sigma^{2},A_{i}}(Y)=\alpha/2.$ (3.6)
The confidence interval
$\mathrm{CI}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}(Y):=\left[a_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)},b_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right]$ (3.7)
has $1-\alpha$ conditional selective coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(Y+\zeta,\delta\right).$
Computing (3.7) requires numerically solving (3.5) for $a_{i}$ and (3.6) for $b_{i}$, where $i=\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$.
Thus, both (3.4) and (3.7) provide confidence intervals for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that attain $1-\alpha$ unconditional selective coverage. We should prefer the approach that yields narrower intervals. Figure 1(a) displays the ratio of the widths of the two intervals when $\mu=0$ and $\sigma^{2}=1$, as we vary the sample size $n$ and the Laplacian variance $2c^{2}$ in (3.3).
The infer-and-widen interval (3.4) involves tuning parameters $\nu$ and $\eta$; to construct Figure 1(a), for a fixed value of the noise $c$ and sample size $n$ in (3.3), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. In the figure, black indicates values of $c$ and $n$ that necessitate a tiny quantile of the normal distribution, yielding numerical issues in R and therefore an interval (3.4) of apparently infinite width. The data fission intervals (3.7) are much narrower than the infer-and-widen intervals (3.4), across all values of $n$ and $c$ considered.
It is possible that the width of the infer-and-widen interval (3.4) is due to the fact that Zrnic and Jordan (2023) apply a series of inequalities in its derivation. Therefore, we now consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $Y_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that achieves a given coverage. It is not computable in practice, as it requires knowledge of the true parameter value $\mu$. In Figures 1(b, c), we numerically compare the “oracle” infer-and-widen method to the data fission interval (3.7) under the selection rule (3.3) with variances $0.332c^{2}=$ 0.33 and $3$, respectively. The former method has a wider average width than the latter: thus, no infer-and-widen interval can outperform the data fission interval (3.4) in this setting.
4 Vignette #2: Maximal contrasts
Following Zrnic and Jordan (2023), we consider the model 3.1 and suppose we also have access to a fixed design matrix with normalized columns $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Our interest lies in the quantity $X_{\tilde{\mathcal{S}}(Y)}^{\top}\mu$ defined via the selection rule
$\tilde{\mathcal{S}}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X_{j}^{\top}Y\right|;$ (4.1)
Zrnic and Jordan (2023) refer to this as the “maximal contrast”. If $X=I$, then this setting is similar to Vignette #1, except with an absolute value in the selection rule.
To achieve a “stable” algorithm, Zrnic and Jordan (2023) consider the selection rule
$\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X^{\top}_{j}Y+\zeta_{j}\right|,$ (4.2)
where $\zeta_{1},\ldots,\zeta_{p}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$ and $c$ is a positive constant that governs the scale of the additive noise.
Proposition 1 (An infer-and-widen interval based on (4.2), given by Zrnic and Jordan (2023)).
Under (3.1), consider the Laplacian selection rule $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ defined in (4.2). For any $\eta>0$ and $\nu\in(0,1)$ such that $c\geq 2z_{1-\alpha\nu/(2p)}/\eta$, the infer-and-widen confidence interval
$\mathrm{CI}_{{{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\nu)e^{-\eta}}(Y):=\left[X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}Y\pm\sigma z_{1-\alpha(1-\nu)e^{-\eta}/2}\right]$ (4.3)
has $1-\alpha$ unconditional selective coverage for $X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\mu$ in the sense of 2.1.
We will take a randomized conditional selective inference approach (Tian and Taylor, 2018; Panigrahi et al., 2024) to construct an alternative confidence interval that makes use of the identical selection event.
Proposition 2 (A randomized conditional selective inference interval based on 4.2).
Consider the model in 3.1 and selection rule $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ in 4.2. Define
$W$ $:=(I_{n}-X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top})Y,$
$\delta$ $:={\rm sign}\left(X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right),$
and $a_{j}$ and $b_{j}$ to satisfy
$\Phi_{a_{j},\sigma^{2},[v_{\min},v_{\max}]}(X_{j}^{\top}Y)=1-\alpha/2$ (4.4)
and
$\Phi_{b_{j},\sigma^{2},[v_{\min},v_{\max}]}(X_{j}^{\top}Y)=\alpha/2,$ (4.5)
where $\Phi_{\lambda,\sigma^{2},[v_{\min},v_{\max}]}(\cdot)$ is defined in Proposition 2, and $v_{\min}$ and $v_{\max}$ are defined in LABEL:sec:deferred. The confidence interval
$\mathrm{CI}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}\left(Y\right):=\left[a_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)},b_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right]$ (4.6)
has $1-\alpha$ conditional selective coverage for $X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y),\zeta,W,\delta\right)$.
Computing 4.6 requires numerically solving 4.4 for $a_{i}$ and 4.5 for $b_{i}$, where $i=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$.
Figure 2: Vignette #2 under (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in 4.2, where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of infer-and-widen 4.3 and randomized conditional selective inference (RCSI) 4.6 intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. (b, c): Fixing $p=100$, we display the average width of the “oracle” infer-and-widen interval and the RCSI interval 4.6 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval 4.3 based on algorithmic stability are also displayed. The Laplacian noise variance is $0.33$ and $3$ in panels (b) and (c), respectively.
We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X\phi$ where half of the elements of $\phi$ are zero and the other half are drawn from an exponential distribution with mean $50/7$. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and its columns are subsequently scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Figure 2(a) displays the ratio of confidence interval widths across values of $p$ and the variance of the Laplacian noise $2c^{2}$ in 4.2.
The infer-and-widen interval (4.6) involves tuning parameters $\nu$ and $\eta$; to construct Figure 2(a), for a fixed value of the noise $c$ and number of features $p$ in (4.2), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. In the figure, black indicates values of $c$ and $p$ that necessitate a tiny quantile of the normal distribution, yielding numerical issues in R and therefore an interval (4.3) of infinite width. The randomized conditional selective inference intervals are narrower than the infer-and-widen intervals across most values of $p$ and $c$ considered.
Once again, we consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top}Y$ that achieves a given coverage. In Figures 2(b, c), we numerically compare the “oracle” infer-and-widen method to the randomized conditional selective inference interval (4.6) under the selection rule (4.2) with variances $0.332c^{2}=$ 0.33 and $3$, respectively. In this setting, the “oracle” infer-and-widen interval is narrower than the randomized conditional selective inference interval in (4.6), thereby indicating that in principle, an infer-and-widen interval can outperform the randomized conditional selective inference interval (4.6) in this setting. However, to the best of our knowledge, no methods to construct such a narrow infer-and-widen interval are currently available.
5 Vignette #3: Inference after the lasso
We now consider inference on the set of coefficients selected by the lasso in a linear model (Tibshirani, 1996). We again consider the model 3.1, alongside a fixed design matrix $X\in\mathbb{R}^{n\times p}$ with unit-normalized columns, i.e., $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Our new selection rule is the magnitude-ordered set of features selected by the lasso:
$\bar{\mathcal{S}}(Y)=$$\bigg{\{}j_{1},\dots,j_{l}:\left|\hat{\beta}_{j_{1}}^{\lambda}\right|>\dots>\left|\hat{\beta}_{j_{l}}^{\lambda}\right|>0,\;$ (5.1)
$\quad\underset{k\notin\{j_{1},\ldots,j_{l}\}}{\max}|\hat{\beta}_{k}^{\lambda}|=0\bigg{\}},$
where
$\hat{\beta}^{\lambda}:=\mathop{\mathrm{arg\,min}}_{\beta}\left\{\left\|Y-X\beta\right\|^{2}_{2}+\lambda\|\beta\|_{1}\right\}.$
We then consider a least squares model fit to the features in $\bar{\mathcal{S}}(Y)$. Our target of inference is the coefficient of the feature with largest lasso magnitude, i.e., $\beta_{{\bar{\mathcal{S}}(Y)},1}$, where
$\beta_{\bar{\mathcal{S}}(Y)}:=A_{\bar{\mathcal{S}}(Y)}\mathbb{E}\left[Y\right]$
and
$A_{J}:=\left(X_{J}^{\top}X_{J}\right)^{-1}X_{J}^{\top}$
for any subset of features $J\subseteq[p]$.
The lasso is not stable in the sense of (Zrnic and Jordan, 2023). Therefore, Zrnic and Jordan (2023) consider a variant of the lasso that is far enough from the original lasso proposal that we do not investigate it here. Instead, in what follows, we construct an interval based on Gaussian data thinning (Neufeld et al., 2024; Dharamshi et al., 2024) that yields conditional guarantees in the sense of 1.3, and then consider the “oracle” infer-and-widen interval that makes use of the identical selection event.
Towards this end, for some $c>0$, we define
$\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)=\bar{\mathcal{S}}(Y+c\zeta),\quad\zeta_{i}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm N}}\,(0,\sigma^{2}).$ (5.2)
Proposition 1.
Under (3.1), consider the selection rule $\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(\cdot)$ defined in (5.2). The data thinning confidence interval
$\mathrm{CI}_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}^{\alpha}(Y):=\left[\hat{\beta}_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y),1}\pm z_{1-\alpha/2}\left(\Sigma_{1,1}\right)^{1/2}\right]$ (5.3)
has $1-\alpha$ conditional selective coverage for $\beta_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y),1}$ in the sense of 1.3 with $\mathcal{C}(Y):=\left(Y+c\zeta\right),$ where
$\hat{\beta}_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}:=A_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}\left(Y-\tfrac{1}{c}\zeta\right)$
and
$\Sigma:=\sigma^{2}\left(1+\tfrac{1}{c^{2}}\right)\left(X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}^{\top}X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}\right)^{-1}.$
We compare the widths of the data thinning (DT) (5.3) and “oracle” infer-and-widen intervals — with identical selection events — in Figure 3. We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X\phi$, where half of the elements of $\phi$ are zero and the other half are either zero or drawn from exponential distributions with means $50/7$ or $5$ in panels (a), (b), and (c) of Figure 3. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and the columns of $X$ are scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. The penalty $\lambda$ is adaptively chosen using 3-fold cross-validation. Figure 3 displays the ratio of the average confidence interval widths as we vary $p$ and the variance of the Gaussian noise $c^{2}$ in (5.2).
We see from Figure 3 that the relative widths of the two intervals depend on the variance of the noise in (5.2) and distribution of $\phi$. Of course, the “oracle” infer-and-widen interval cannot be computed in practice: any real-world infer-and-widen interval that achieves nominal coverage would necessarily be at least as wide as the oracle interval, if not far wider. Thus, while it may be possible to develop an infer-and-widen interval that might outperform the simple data thinning interval in 5.3 in specific scenarios, the potential upside of any such approach is extremely limited.
Figure 3: Vignette #3 under (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in 5.2. The columns of $X$ are independent normal draws with column correlation $0.5$. Results are averaged over $1000$ simulated datasets. We consider inference on a lasso-selected coefficient in a least squares model, varying the mean of $\phi$ across $0,5,50/7$ in panels (a), (b), and (c), respectively. We display the ratio of average widths of the 95% “oracle” infer-and-widen interval and the 95% data thinning (DT) interval (5.3), with identical selection events and level $\alpha=0.05$; a positive log ratio indicates that the data thinning interval is narrower and hence preferable. In many settings, even the unattainable “oracle” infer-and-widen interval is wider than the attainable data thinning interval.
6 The bias of infer-and-widen intervals
We now investigate why infer-and-widen intervals are too wide. To facilitate the analysis, we consider Vignette #1 with a simple selection rule that uses Gaussian noise,
$\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}(Y_{i}+c\zeta_{i}),\;\;\;\;\zeta_{i}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm N}}\,(0,\sigma^{2}).$ (6.1)
We know from Section 2 that any infer-and-widen interval is centered around $Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}$.
Proposition 1 (Bounding the bias of the infer-and-widen midpoint).
Under (3.1), with $\mu_{1}=\dots=\mu_{n}$, consider the selection rule $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)$ in (6.1). Then
$\mathbb{E}\left[Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]\geq\frac{\sigma}{2}(\sqrt{1+c^{2}}-c)\sqrt{\log n}.$
That is, the bias increases on the order of $\sqrt{\log n}$. Note that as $c\to\infty$, i.e., as the selection rule becomes fully random, this lower bound goes to zero as expected.
Noting that $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)$ in (6.1) is independent of $Y-\frac{1}{c}\zeta$ (Neufeld et al., 2024), it is straightforward to show that $[(Y-\frac{1}{c}\zeta)_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\pm\sigma(1+c^{-2})^{1/2}z_{1-\alpha/2}]$ is a valid $1-\alpha$ selective confidence interval for $\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}$. Furthermore,
$\mathbb{E}\left[\left(Y-\frac{1}{c}\zeta\right)_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]=0,$
i.e., the midpoint is unbiased for the mean of the selected observation.
Thus, infer-and-widen here leads to a confidence interval that is centered in the wrong place, but a simple alternative that makes use of the same selection event yields an interval that is correctly centered.
7 Discussion
In this paper, we compared the performance of the “infer-and-widen” approach to simple alternative methods in three vignettes: two taken from Zrnic and Jordan (2023), and the third involving feature selection via the lasso. Our findings are as follows:
  1. 1.
    In the first two vignettes, infer-and-widen intervals can be computed in practice but are substantially wider than intervals constructed using alternative approaches that make use of identical randomized selection events and yield conditional guarantees.
  2. 2.
    In general, for a given nominal coverage, the “oracle” infer-and-widen interval is substantially narrower than an infer-and-widen interval that can be computed in practice.
  3. 3.
    In Vignettes #1 and #3, the “oracle” infer-and-widen interval was often wider than the alternative approaches we consider; thus, there may be limited value in developing narrower infer-and-widen methods that can be computed in practice.
One may wonder whether our findings conflict with those of Goeman and Solari (2022), who proved that any method that guarantees conditional coverage (1.3) is dominated by an unconditional multiple testing procedure defined on the full universe of hypotheses. However, their setting is quite different from ours: they consider inference on non-selected parameters and unconditional procedures outside of the infer-and-widen framework.
Acknowledgements
We acknowledge funding from the following sources: ONR N00014-23-1-2589, NSF DMS 2322920, a Simons Investigator Award in Mathematical Modeling of Living Systems, and NIH 5P30DA048736.
Appendix A A summary of algorithmic stability
The notion of stability underlying the algorithmic stability proposal of [Zrnic and Jordan, 2023] is as follows.
Definition 1 (Definitions 4.1 and 4.2, Zrnic and Jordan [2023]).
Two random variables $Q$ and $T$ are $(\eta,\tau)$-indistinguishable from each other, denoted by $Q\approx_{(\eta,\tau)}T$, if for any measurable set $\mathcal{O}$, it holds that
$\mathbb{P}(Q\in\mathcal{O})\leq e^{\eta}\mathbb{P}(T\in\mathcal{O})+\tau.$ (A.1)
A (possibly randomized) selection rule ${\mathcal{S}}:\mathcal{Y}\mapsto\mathcal{I}$ is $(\eta,\tau,\nu)$-stable with respect to a distribution $\mathcal{P}$ on $\mathcal{Y}$ if there exists a (possibly random, possibly dependent on $\mathcal{P}$) variable ${S}_{0}$ such that
$\mathbb{P}\left(\mathcal{S}(Y)\approx_{(\eta,\tau)}{S}_{0}\right)\geq 1-\nu.$ (A.2)
Note that the probability is taken solely with respect to $Y\sim\mathcal{P}$.
Since $S_{0}$ is independent of $Y$, we can use the classical interval (1.1) to conduct inference on a target given by $S_{0}$, i.e., it holds that $\mathbb{P}\left(\theta_{S_{0}}\not\in\mathrm{CI}_{S_{0}}^{\alpha}(Y)\right)\leq\alpha$. The key insight underlying algorithmic stability is that if $\mathcal{S}(Y)$ is stable, then for most $Y\sim\mathcal{P}$ the value of $\mathcal{S}(Y)$ is close (in the sense of Definition 1) to $S_{0}$. Thus, it is “almost as if” the selection rule $\mathcal{S}(Y)$ did not involve $Y$. This notion is formalized in Theorem 2.
Theorem 2 (Theorem 5.2, Zrnic and Jordan [2023]).
Let $\mathcal{S}(\cdot)$ be a $(\eta,\tau,\nu)$-stable selection rule with parameters possibly dependent on the data distribution $\mathcal{P}$ of $Y$. Assume that for any fixed $\theta_{i}\in\Theta$, we have access to a confidence interval $\mathrm{CI}_{i}^{\alpha}(Y)$ for $\theta_{i}$ with coverage at least $1-\alpha$, i.e., it satisfies (1.1). Then
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{(\alpha-\tau-\nu)e^{-\eta}}(Y)\right)\leq\alpha.$ (A.3)
Thus, to conduct inference on a data-driven target $\theta_{\mathcal{S}(Y)}$, we can apply the infer-and-widen interval in (2.1) with $H(\alpha,{\mathcal{S}})=(\alpha-\tau-\nu)e^{-\eta}$, a function of the stability of the selection rule $\mathcal{S}(\cdot)$. The values of $\eta,\tau,$ and $\nu$ that govern the stability of $\mathcal{S}(\cdot)$ directly determine the width of the confidence interval: the smaller these parameters, the narrower the confidence interval. Zrnic and Jordan [2023] construct stable selection rules in a number of settings via the introduction of auxiliary randomness.
Appendix B Vignette #1
B.1 Proof of Proposition 2
To construct a confidence interval with coverage as in (1.3) under the selection rule (3.2), it suffices to characterize the distribution $Y\mid Y^{\mathrm{train}}$, where $Y^{\mathrm{train}}:=Y+\zeta$ and $\zeta_{1},\ldots,\zeta_{n}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$. In the language of Leiner et al. [2023], this is the “data fission” approach. This conditional distribution is difficult to characterize, however, and so we instead consider the distribution conditional on $Y^{\mathrm{train}}$ and an additional event.
Lemma 1 (Data fission with Laplace noise).
Let $Y\sim{\bf{\rm N}}\,_{n}(\mu,\sigma^{2}I_{n})$ where $\mu\in\mathbb{R}^{n}$. Let $\zeta\in\mathbb{R}^{n}$ where $\zeta_{i}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm Laplace}}\,(c)$. Define $Y^{\mathrm{train}}:=Y+\zeta$, $\delta:={\rm sign}(Y^{\mathrm{train}}-Y)$, and the set $A:=\left\{y\in\mathbb{R}^{n}:\delta_{i}y_{i}\leq\delta_{i}Y^{\mathrm{train}}_{i}\;\forall i\in[n]\right\}$. Then the density of the conditional distribution $Y_{i}\mid(Y^{\mathrm{train}},\delta)$ has the following form:
$Y\mid(Y^{\mathrm{train}},\delta)\sim{\bf{\rm N}}\,_{n,A}(\mu+c^{-1}\sigma^{2}\delta,\sigma^{2}I_{n}),$ (B.1)
where ${\bf{\rm N}}\,_{n,A}$ denotes an $n$-dimensional multivariate normal distribution truncated to the set $A$.
Proof.
The conditional density is as follows:
$f_{Y|(Y^{\mathrm{train}},\delta)}(Y)$ $\propto f_{\delta|Y,Y^{\mathrm{train}}}(\delta)f_{Y^{\mathrm{train}}|Y}(Y^{\mathrm{train}})f_{Y}(Y)$
$\propto\mathbb{I}[Y\in A]f_{Y^{\mathrm{train}}-Y|Y}(Y^{\mathrm{train}}-Y)f_{Y}(Y)$
$=\mathbb{I}[Y\in A]f_{\zeta}(Y^{\mathrm{train}}-Y)f_{Y}(Y)$
$=\mathbb{I}[Y\in A]\prod_{i=1}^{n}f_{\zeta_{i}}(Y^{\mathrm{train}}_{i}-Y_{i})f_{Y}(Y)$
$\propto\mathbb{I}[Y\in A]\exp\left\{-\frac{1}{2\sigma^{2}}(Y-\mu)^{\top}(Y-\mu)-c^{-1}\sum_{i=1}^{n}|Y^{\mathrm{train}}_{i}-Y_{i}|\right\}$
$=\mathbb{I}[Y\in A]\exp\left\{-\frac{1}{2\sigma^{2}}(Y-\mu)^{\top}(Y-\mu)+c^{-1}\delta^{\top}(Y-Y^{\mathrm{train}})\right\}$
$\propto\mathbb{I}[Y\in A]\exp\left\{-\frac{1}{2\sigma^{2}}(Y-\mu-c^{-1}\sigma^{2}\delta)^{\top}(Y-\mu-c^{-1}\sigma^{2}\delta)\right\}.$ (B.2)
Thus the density of $Y|(Y^{\mathrm{train}},\delta)$ is that of multivariate normal distribution with independent marginals, truncated to $A$, which restricts the domain to the intersection of halfspaces defined by elements of $Y^{\mathrm{train}}$. ∎
It remains to derive a confidence interval for $\mu_{i}$. Note that since $Y\mid(Y^{\mathrm{train}},\delta)$ is a truncated normal distribution with a diagonal covariance, the marginal distributions are also truncated normal distributions (note that this is not in general true for an arbitrary covariance). Thus the marginal distribution $Y_{i}\mid(Y^{\mathrm{train}},\delta)\sim{\bf{\rm N}}\,_{A_{i}}\left(\mu_{i}+c^{-1}\sigma^{2}\delta_{i},\sigma^{2}\right)$, where $A_{i}:=\left\{y\in\mathbb{R}^{n}:\delta_{i}y_{i}\leq\delta_{i}Y^{\mathrm{train}}_{i}\right\}.$ A valid post-selection confidence interval for $\mu_{i}$ follows from the CDF of this truncated normal distribution, yielding Proposition 2.
B.2 Proof of Proposition 6
First, without loss of generality, let $\mu_{1}=\dots=\mu_{n}=0$. Then
$\mathbb{E}\left[Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]=\mathbb{E}\left[Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]=\mathbb{E}\left[\mathbb{E}\left[Y_{i}|i=\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)\right]\right].$
To bound the conditional expectation, let $X_{i}:=Y_{i}+\zeta_{i}$, where $\zeta_{i}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm N}}\,(0,c^{2}\sigma^{2})$ is the additive noise. Thus, $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)=\mathcal{S}(X)$. Let $Z\sim{\bf{\rm N}}\,_{n}(0,I_{n})$, with maximum order statistic $Z_{(1)}=Z_{\mathcal{S}(Z)}$. We can bound the conditional expectation
$\mathbb{E}\left[Y_{i}|i=\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)\right]$ $=\mathbb{E}\left[X_{i}-\zeta_{i}|i=\mathcal{S}(X)\right]$
$=\mathbb{E}\left[X_{i}|i=\mathcal{S}(X)\right]-\mathbb{E}\left[\zeta_{i}|i=\mathcal{S}(X)\right]$
$\geq\mathbb{E}\left[X_{i}|i=\mathcal{S}(X)\right]-\mathbb{E}\left[\zeta_{i}|i=\mathcal{S}(\zeta)\right]$
$=\sigma\sqrt{1+c^{2}}\mathbb{E}\left[Z_{i}|i=\mathcal{S}(Z)\right]-\sigma c\mathbb{E}\left[Z_{i}|i=\mathcal{S}(Z)\right]$
$=\sigma(\sqrt{1+c^{2}}-c)\mathbb{E}[Z_{(1)}]$
$\geq\frac{\sigma}{2}(\sqrt{1+c^{2}}-c)\sqrt{\log n},$
where the final inequality comes from Kamath [2015]. Taking the expectation on each side completes the proof.
Appendix C Vignette #2
C.1 Proof of Proposition 4
Proof.
The proof will make use of Theorem 5.2 of Lee et al. [2016], which we restate now.
Theorem 1 (Theorem 5.2 of Lee et al. [2016]).
Let $Y\sim{\bf{\rm N}}\,_{n}(\mu,\Sigma)$ be an $n$-dimensional random vector, and $\beta\in\mathbb{R}^{n}$ a vector of interest. For any fixed matrix $D\in\mathbb{R}^{n\times d}$ and vector $b\in\mathbb{R}^{d}$, define
$\gamma:=\Sigma\beta(\beta^{\top}\Sigma\beta)^{-1}\quad\mathrm{and}\quad W:=(I_{n}-\gamma\beta^{\top})Y,$
and the pair
$v_{\min}:=\underset{j:(D\gamma)_{j}<0}{\mathrm{max}}(b_{j}-(DW)_{j})/(D\gamma)_{j}\quad\mathrm{and}\quad v_{\max}:=\min_{j:(D\gamma)_{j}>0}(b_{j}-(DW)_{j})/(D\gamma)_{j}.$
Then
$\beta^{\top}Y|\left(DY\leq b,W\right)\sim{\bf{\rm N}}\,_{[v_{\min},v_{\max}]}(\beta^{\top}\mu,\beta^{\top}\Sigma\beta),$ (C.1)
where ${\bf{\rm N}}\,_{[v_{\min},v_{\max}]}$ denotes a normal distribution truncated to $[v_{\min},v_{\max}]$.
Returning to our setting in Proposition 4, for brevity let $Z:=X^{\top}Y$ and $Z^{\mathrm{train}}:=Z+\zeta$ where $\zeta_{1},\ldots,\zeta_{p}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm Laplace}}\,(c)$. Recall that our target of inference is $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$, where $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ is the selection rule in (4.2). This selection rule is implicitly a function of the Laplacian noise $\zeta$, and so to emphasize this point we will write $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)$ in this section.
To construct a confidence interval with conditional coverage as in (1.3), it will suffice to derive the distribution of $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y\mid C(Y,\zeta)$ for an event $\mathcal{C}(Y)$ satisfying $\sigma\left(C(Y,\zeta)\right)\supseteq\sigma\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)\right)$, where $\sigma(\cdot)$ denotes the sigma algebra generated by a random variable, e.g., we will condition on the selection event along with an additional event.
We will show that there exists a matrix $A_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\in\mathbb{R}^{(2p+1)\times p}$ for which
$\left\{(Y,\zeta):\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})\right\}=\left\{(Y,\zeta):\delta A_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}X^{\top}Y\leq-\delta A_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\zeta\right\},$ (C.2)
where
$\delta:={\rm sign}\left(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right)={\rm sign}\left(Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right).$ (C.3)
Henceforth, for brevity, we will denote $A_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}$ by $A$.
To show that there exists a matrix $A$ for which (C.2) holds, note that by definition $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $|Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}|\geq|Z^{\mathrm{train}}_{j}|$ for all $j\in[p]$. In turn, expanding the absolute value and definition of $Z^{\mathrm{train}}$, (C.2) holds if and only if both (I) and (II) hold:
  1. (I)
    if $Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\geq 0$, then for all $j\neq\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ it holds that
    1. (a)
      $Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\geq 0$,
    2. (b)
      $Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\geq Z_{j}+\zeta_{j}$,
    3. (c)
      $Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\geq-(Z_{j}+\zeta_{j})$;
  2. (II)
    else if $Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq 0$, then for all $j\neq\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ it holds that
    1. (a)
      $Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq 0$,
    2. (b)
      $Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq Z_{j}+\zeta_{j}$,
    3. (c)
      $Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq-(Z_{j}+\zeta_{j})$.
We further rearrange the above inequalities to collect the $Z$ terms on one side, as follows:
  1. (I)
    If $Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\geq 0$, then for all $j\neq\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ it holds that
    1. (a)
      $-Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}$,
    2. (b)
      $Z_{j}-Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}-\zeta_{j}$,
    3. (c)
      $-Z_{j}-Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}+\zeta_{j}$;
  2. (II)
    else if $Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq 0$, then for all $j\neq\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ it holds that
    1. (a)
      $Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq-\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}$,
    2. (b)
      $Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}-Z_{j}\leq\zeta_{j}-\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}$,
    3. (c)
      $Z_{j}+Z_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq-\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}-\zeta_{j}$.
Thus, if $Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\geq 0$, then we are in case (I), and $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $Z$ satisfies $AZ\leq-A\zeta$, where
$A:=\begin{bmatrix}0_{1\times n_{1}}&-1&0_{1\times n_{2}}\\ I_{n_{1}}&-1_{n_{1}}&0_{n_{1}\times n_{2}}\\ 0_{n_{2}\times n_{1}}&-1_{n_{2}}&I_{n_{2}}\\ -I_{n_{1}}&-1_{n_{1}}&0_{n_{1}\times n_{2}}\\ 0_{n_{2}\times n_{1}}&-1_{n_{2}}&-I_{n_{2}}\end{bmatrix},$ (C.4)
and we denote $n_{1}:=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})-1$ and $n_{2}:=p-\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ for brevity. If instead $Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq 0$, then we are in case (II), where $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $Z$ satisfies $-AZ\leq A\zeta$. Thus, in either case, $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $\delta AZ\leq-\delta A\zeta$, and so, by definition of $Z$, if and only if $\delta AX^{\top}Y\leq-\delta A\zeta$. This establishes that (C.2) holds for $A$ defined in (C.4).
Thus, by C.2, we see that
$X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})\right)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y|\left(\delta AX^{\top}Y\leq-\delta A\zeta\right).$
We would like to apply Theorem 1 to the right-hand side of this result with $D=\delta AX^{\top}$ and $b=-\delta A\zeta$, but these are random quantities and so we will need to fix them by conditioning on $\zeta$ and $\delta$.
Defining $W=(I_{n}-X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top})Y$ as in Theorem 1, we have that
$X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs}),\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)$ (C.5)
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs}),\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)$
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta AX^{\top}Y\leq-\delta A\zeta,\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)$ (C.6)
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta,\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)$
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta,\zeta=\zeta^{obs},W\right)$ (C.7)
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta^{obs},\zeta=\zeta^{obs},W\right)$
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta^{obs},W\right).$ (C.8)
C.6 follows from the equivalence in C.2.
To establish C.7, note that the event $\{\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta\}$ is composed of a system of $12p+$ 1 inequalities, the first of which states that $-\delta^{obs}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y\leq\delta^{obs}\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}$, or equivalently, that $0\leq\delta^{obs}\left(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right)$. Recalling the definition of $\delta$ in (C.3), it follows that the event $\{\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta\}$ implies that $\delta=\delta^{obs}$. This yields (C.7).
[…]
We now apply Theorem 1 with $Y\sim{\bf{\rm N}}\,_{n}(\mu,\sigma^{2}I_{n})$, $\beta=X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}$, $D=\delta^{obs}AX^{\top}$, and $b=-\delta^{obs}A\zeta^{obs}$. Furthermore, recalling that $X$ is column-normalized, we have that $\gamma=X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta^{obs})}$ and $W=\left(I_{n}-X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}\right)Y$ in the statement of Theorem 1. Thus,
$X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta^{obs},W\right)\sim{\bf{\rm N}}\,_{[v_{\min},v_{\max}]}(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}\mu,\sigma^{2}),$ (C.9)
for $v_{\min}$ and $v_{\max}$ defined by
$v_{j}$ $:=-(A\zeta+AX^{\top}W)_{j}/(AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)})_{j},$
$v_{\max}$ $:=\underset{j:\left(\delta AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right)_{j}<0}{\mathrm{max}}v_{j},$
$v_{\min}$ $:=\min_{j:\left(\delta AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right)_{j}>0}v_{j}.$
Applying C.5-C.8 to (C.9), we have that
$X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs}),\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)\sim{\bf{\rm N}}\,_{[v_{\min},v_{\max}]}(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}\mu,\sigma^{2}).$
A confidence interval for $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}\mu~{}|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta),\zeta,\delta,W\right)$ with conditional coverage guarantee (1.3), where $\mathcal{C}(Y)=\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y),\zeta,\delta,W\right)$, is thus obtained using the CDF of the above truncated normal distribution. ∎
C.2 Additional simulation results
The simulation results in Section 4 reflect a limited set of parameter choices. Figure A1 extends Figure 2(a) to a wider range of levels of signal. Figure A2 extends Figure 2(b, c) to a wider range of levels of signal.
(a) $\phi=0$                                (b) $\mathbb{E}[\phi]=5$                            (c) $\mathbb{E}[\phi]=\tfrac{50}{7}$ Refer to caption
Figure A1: Vignette #2, under the model (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in (4.2), where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. We vary the mean of $\phi$ across $0,5,50/7$ in panels (a), (b), and (c), respectively. Results are averaged over $250$ simulated datasets. We display the ratio of mean widths of infer-and-widen (4.3) and randomized conditional selective inference (RCSI) (4.6) intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the average infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. The differences in widths is most pronounced under the null in panel (a), but results are qualitatively similar across levels of signal.
$\phi=0$                               $\mathbb{E}[\phi]=5$                               $\mathbb{E}[\phi]=\tfrac{50}{7}$ Refer to caption
Figure A2: Vignette #2, under the model (3.1) with $n=100$, $p=100$, $\sigma=1$, $\mu=\phi$, and the selection rule in (4.2), where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. We vary the mean of $\phi$ across the three columns, and the Laplacian noise variance across the two rows. We display the average width of the “oracle” infer-and-widen interval and the randomized conditional selective inference (RCSI) interval (4.6) as a function of coverage. The 95% classical interval, and the “attainable” 95% infer-and-widen interval (4.3) based on algorithmic stability, are also displayed. Results are qualitatively similar regardless of choice of parameters.
Appendix D Vignette #3
D.1 Proof of Proposition 5
The result follows from independence of the selection and inference stages, per the following result.
Lemma 1 (Gaussian data thinning, Rasines and Young [2023], Neufeld et al. [2024]).
[…]
The selection event $\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)$ in (5.3) is a function of $Y$ through $Y^{\mathrm{train}}$ and is independent of $Y^{\mathrm{test}}$. We proceed to conduct conventional inference using $Y^{\mathrm{test}}$ in an ordinary least squares model on the coefficient selected using $Y^{\mathrm{train}}$, as given in Section 4.1 of  Berk et al. [2013]. This yields a conditional guarantee of the form in (1.3), with $\mathcal{C}(Y)=(Y^{\mathrm{train}}).$