### v3

# Infer-and-widen, or not?

Abstract
In recent years, there has been substantial interest in the task of selective inference: inference on a parameter that is selected from the data. Many of the existing proposals fall into what we refer to as the infer-and-widen framework: they produce symmetric confidence intervals whose midpoints do not account for selection and therefore are biased; thus, the intervals must be wide enough to account for this bias. In this paper, we investigate infer-and-widen approaches in three vignettes: the winner’s curse, maximal contrasts, and inference after the lasso. In each of these examples, we show that a state-of-the-art infer-and-widen proposal leads to confidence intervals that are wider than a non-infer-and-widen alternative. Furthermore, even an “oracle” infer-and-widen confidence interval — the narrowest possible interval that could be theoretically attained via infer-and-widen — can be wider than the alternative.
infer-and-widen,
keywords:
[…]
, , , and
1 Introduction
In classical statistics, the target of inference — for instance, the null hypothesis that we wish to test, or the parameter for which we wish to construct a confidence interval — is fixed. However, motivated by the explosion of data across a variety of fields, recent interest has focused on the setting where the target of inference is itself a function of the data, and therefore is random. The terms selective inference and post-selection inference describe the set of strategies available for conducting inference on a random target.
To formalize this, consider a random variable $Y$ with domain $\mathcal{Y}$, and a collection of parameters $\theta$ whose elements can be indexed by a (possibly infinite) index set $\mathcal{I}$. For a fixed index $i\in\mathcal{I}$, and for $\alpha\in(0,1)$, a classical $1-\alpha$ confidence interval for $\theta_{i}$, which we will denote $\mathrm{CI}_{i}^{\alpha}$, satisfies
$\mathbb{P}\left(\theta_{i}\not\in\mathrm{CI}_{i}^{\alpha}(Y)\right)\leq\alpha.$ (1.1)
Now, suppose that $i\in\mathcal{I}$ is no longer a fixed index; instead, some selection rule ${\mathcal{S}}:\mathcal{Y}\to\mathcal{I}$ uses the data to select a parameter on which to conduct inference. A confidence interval provides valid unconditional coverage [3] at level $\alpha$ if
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{\alpha}(Y)\right)\leq\alpha.$ (1.2)
Unconditional coverage is not guaranteed by classical confidence intervals. In this paper, we consider methods that yield unconditional guarantees, in the sense of 1.2. Some of these confidence intervals also yield (stronger) conditional guarantees [10], in the sense that
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{\alpha}(Y)~|~\mathcal{C}(Y)\right)\leq\alpha$ (1.3)
for an event $\mathcal{C}(Y)$ satisfying $\sigma\left(\mathcal{C}(Y)\right)\supseteq\sigma\left({\mathcal{S}}(Y)\right)$, where $\sigma(\cdot)$ denotes the sigma algebra generated by a random variable, i.e., the coverage guarantee in 1.3 conditions on at least the selection event. By the law of total expectation, conditional coverage, 1.3, implies unconditional coverage, 1.2.
1.1 The infer-and-widen framework
As presented in Box 1.1, the infer-and-widen framework involves (Step 1) defining a target of inference using some or all of the data, and then (Step 2) symmetrically adjusting the critical values arising from “classical” inference to account for the fact that the target of inference is not fixed in order to achieve unconditional coverage, (1.2). The simplest examples of this strategy are the Bonferroni [4, 8], Holm-Bonferroni [14], and Šidák [23] procedures, which are suitable when a finite number of targets of inference are considered (and in the case of Šidák, under additional assumptions). In the special case of inference on contrasts of a linear model, the Scheffé correction is available [22]; this idea is further refined by the PoSI method [3]. While conceptually attractive, these approaches tend to be quite conservative (Bonferroni, Holm-Bonferroni, Šidák), are limited to a linear model setting (Scheffé, PoSI), or are often computationally intractable (PoSI).
Other recent infer-and-widen approaches have attempted to address these shortcomings. The local simultaneous inference approach of Zrnic and Fithian [26] focuses a simultaneous correction on the subset of hypotheses that are most likely given the observed data. The algorithmic stability proposal of [28] also adapts to the details of the selection algorithm. By drawing connections with the field of differential privacy [9], the “stability” of the approach for selecting the target of inference in Step 1 of Box 1.1 is quantified; this, in turn, leads to a stability-specific adjustment of the critical values in Step 2 of Box 1.1 (see also Section A of the online supplement). In a sense, these two approaches achieve the lofty goal laid out in the PoSI proposal of [3] by creating frameworks to compute selection-specific adjustments to the critical values.
Box 1: Infer-and-widen. 1. A target of inference is selected from the data. 2. “Classical” inference is conducted, with a symmetric adjustment of the critical values to account for the data-driven selection of the target of inference.
1.2 Other approaches for valid selective inference
Not all approaches that yield unconditional guarantees of the form (1.2) fall into the infer-and-widen framework. Some methods produce asymmetric confidence intervals around the biased point estimate: for instance, Benjamini et al. [2] and Fuentes et al. [11] construct an asymmetric Bonferroni correction that accounts for the direction of the bias. More recently, Zrnic and Fithian [27] extend the local simultaneous approach of Zrnic and Fithian [26] by additionally adapting the correction to the specific selection algorithm. The hybrid approaches of [1] and [18] apportion a fraction of the Type 1 error budget to rule out unlikely selection events, and then construct a conditionally-valid confidence interval on the likely hypotheses; note that this only guarantees unconditional coverage, 1.2. Lastly, there is a line of empirical Bayes work whose intervals are centered at shrinkage estimators which, while likely biased, improve upon the non-selective estimator [15].
Furthermore, a number of methods yield coverage guarantees conditional on the selection event, i.e., (1.3). Some methods do this by obtaining independent training and test sets; then, the training set can be used to define the target of inference, and the test set to conduct inference on this target. Because the two sets are independent, we can conduct inference using the test set as though the target of inference were fixed, despite the fact that it is actually a function of the training data. Sample splitting, applicable in the setting of independent and identically distributed observations, provides one approach to obtain independent training and test sets [6]. More recent data thinning proposals generalize the idea of sample splitting, and can be applied when only a single observation is available [21, 19, 7].
In cases where it is either impossible or undesirable to create independent training and test sets, strategies that achieve conditional guarantees of the form 1.3 may still be available. For instance, a target of inference can be defined using all of the data, and then inference can be conducted conditional on the event that the data led to this particular target of inference. Examples include the proposals of [10, 16, 5, 12, 1]. Alternatively, if a researcher is willing to define the target of inference using only a subset of the data or using a noisy version of the data, then data fission [17] or randomized conditional selective inference [24, 20] may be applicable. Randomization in this manner preserves information for inference, overcoming a limitation of conditioning on selection using all of the data.
1.3 Contributions and organization
In this paper, we conduct an empirical and theoretical evaluation of infer-and-widen strategies in the context of three vignettes. The first two vignettes involve inference after a randomized selection rule, and we derive alternative confidence intervals to compare to an existing infer-and-widen approach. In the third vignette, we compare a suite of existing methods. The bottom line of our paper is as follows: when comparing infer-and-widen strategies to alternative selective inference methods that make use of identical selection events, the latter often outperform the former in terms of confidence interval width. Furthermore, in some settings, the latter even outperform the “oracle” infer-and-widen procedure: the narrowest possible interval that could be theoretically attained within the infer-and-widen framework.
The organization is as follows. In Section 2, we formalize the infer-and-widen framework. In Sections 3, 4, and 5, we compare its performance to new alternative intervals that we derive in two vignettes considered in [28], and to existing alternative intervals in an example involving model selection via the lasso. In Sections 3 and 4, the selection event is randomized, whereas in Section 5 it is not. In Section 6, we show analytically that the infer-and-widen interval can suffer from a highly biased midpoint. We close with a discussion in Section 7. Theoretical results and additional empirical comparisons are proven in the online supplement. Scripts to fully reproduce all numerical results can be found at https://github.com/rflperry/infer_and_widen.
2 Formalizing the infer-and-widen framework
We now formalize the infer-and-widen (I&W) framework, outlined in Box 1.1. This involves a modification of the classical confidence interval $\mathrm{CI}_{i}^{\alpha}$ in (1.1) to account for the fact that $\theta_{\mathcal{S}(Y)}$ is a data-driven parameter. Based on the particulars of the selection function $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ and data $Y$, a modified $\alpha$-level, $H(\alpha,{\mathcal{S}},Y)$, is derived so that $\mathrm{CI}^{H(\alpha,{\mathcal{S}},Y)}_{\mathcal{S}(Y)}(Y)$ is unconditionally valid at level $\alpha$ in the sense of 1.2, i.e.,
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{H(\alpha,{\mathcal{S}},Y)}(Y)\right)\leq\alpha.$ (2.1)
In effect, this procedure symmetrically widens the classical confidence interval to account for bias in its midpoint.
In a Bonferroni correction we have $H(\alpha,{\mathcal{S}},Y)=\alpha/|\mathcal{I}|$, and in the Holm-Bonferroni correction it is slightly larger. Since these approaches adapt neither to the selection rule $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ nor to the data $Y$, the constant $H(\alpha,{\mathcal{S}},Y)$ is easy to compute but overly conservative. Despite the notation $H(\alpha,{\mathcal{S}},Y)$ used in (2.1), the aforementioned infer-and-widen proposals result in confidence intervals whose widths do not depend on the particulars of the selection rule $\mathcal{S}(\cdot)$ nor data $Y$.
The foundational PoSI proposal [3] considers the setting of linear regression with fixed regressors. They derive a constant $H(\alpha,{\mathcal{S}},Y)$ that enables inference on the coefficient(s) in any submodel, regardless of the form of $\mathcal{S}(\cdot)$ or observed data $Y$. However, for a particular selection rule $\mathcal{S}(\cdot)$, the resulting confidence interval may be quite conservative, and computing the constant $H(\alpha,{\mathcal{S}},Y)$ is challenging or infeasible.
In contrast are two newer approaches, which provide narrower constants $H(\alpha,{\mathcal{S}},Y)$ that depend on the selection rule ${\mathcal{S}}(\cdot)$. The locally simultaneous inference method of Zrnic and Fithian [26] apportions the Type-1 error budget $\alpha$, first to use the data $Y$ to rule out the most unlikely hypotheses, and then to perform a simultaneous correction over a subset of all $\left|\mathcal{I}\right|$ hypotheses. The algorithmic stability proposal of [28] adjusts the critical value $H(\alpha,{\mathcal{S}},Y)$ based on how “stable” the selection rule ${\mathcal{S}}(\cdot)$ is; intuitively, a more “stable” selection rule requires a less substantial adjustment. Section A in the online supplement contains a more detailed overview.
We close with a remark about the infer-and-widen framework: at first glance, it appears that an “advantage” of infer-and-widen, over alternative selective inference approaches, is that the interval in (2.1) is of the same form as the classical interval in the absence of selection (1.1), i.e., no bespoke methods for computing a confidence interval are required. However, this is a double-edged sword, in the sense that bias induced by the selection event is perpetuated into the confidence interval. Thus, an infer-and-widen confidence interval that attains the nominal coverage may be extremely wide, to account for bias in the interval’s midpoint. We will see that this is a major drawback of infer-and-widen.
3 Vignette #1: The winner’s curse
Suppose that we observe $n$ independent Gaussian random variables $Y_{1},\ldots,Y_{n}$ with common known variance, and are interested in conducting inference on the unknown mean of the largest observed variable. That is,
$Y\sim{\bf{\rm N}}\,_{n}(\mu,\sigma^{2}I_{n}),$ (3.1)
and the selection rule is
$\mathcal{S}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}Y_{i}.$ (3.2)
It is not hard to see that a $1-\alpha$ confidence interval for $\mu_{\mathcal{S}(Y)}$ that treats $\mathcal{S}(Y)$ as though it were fixed (rather than a function of the data) will fail to achieve the nominal coverage. This is the “winner’s curse” [1].
To achieve a “stable” algorithm, [28] consider a randomized version of the selection rule (3.2) via the addition of centered Laplace noise,
$\mathcal{S}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}(Y_{i}+\zeta_{i}),$ (3.3)
where $\zeta_{1},\ldots,\zeta_{n}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$ and $c$ is a positive constant that governs the amount of additive noise. When $c=0$, $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ simplifies to $\mathcal{S}(Y)$ in (3.2).
Proposition 1 (An infer-and-widen interval based on (3.3), given by [28]).
Under 3.1, consider the selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(\cdot)$ defined in (3.3). For any $\eta>0$ and $\nu\in(0,\alpha)$ such that $\eta c\geq 2z_{1-\alpha(\alpha-\nu)/2n}$, where $z_{q}$ denotes the $q$th quantile of the standard normal distribution, the infer-and-widen confidence interval
$\mathrm{CI}_{{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\alpha+\nu)e^{-\eta}}(Y):=\left[Y_{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}\pm\sigma z_{1-\alpha(1-\alpha+\nu)e^{-\eta}/2}\right]$ (3.4)
has $1-\alpha$ unconditional coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 2.1.
In the context of Proposition 1, the selection algorithm in 3.3 is “stable” at level $\eta$ with probability at least $1-\nu$. We defer a more detailed discussion to Section A in the online supplement.
We will now derive an alternative interval that makes use of exactly the same selection event, (3.3), using ideas from the data fission proposal of [17]. Towards this end, note that $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)=\mathcal{S}(Y^{\mathrm{train}})$ where $Y^{\mathrm{train}}:=Y+\zeta$. Since $Y\mid Y^{\mathrm{train}}$ does not admit a tractable distribution, we will additionally condition on $\left({\rm sign}(\zeta_{1}),\ldots,{\rm sign}(\zeta_{n})\right)$. This yields a multivariate truncated normal distribution with independent marginals.
Proposition 2 (A data fission interval based on (3.3)).
Under (3.1), consider the selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ defined in (3.3). Define $\delta_{i}:={\rm sign}(\zeta_{i})$ and $A_{i}:=\{y_{i}\in\mathbb{R}:\delta_{i}y_{i}\leq\delta_{i}(Y_{i}+\zeta_{i})\}$. Let $\Phi_{\mu,\sigma^{2},A_{i}}(\cdot)$ denote the CDF of a ${\bf{\rm N}}\,(\mu,\sigma^{2})$ distribution truncated to the set $A_{i}$, and define $a_{i}$ and $b_{i}$ to satisfy
$\Phi_{a_{i}+\delta_{i}\sigma^{2}/c,\sigma^{2},A_{i}}(Y)=1-\alpha/2$ (3.5)
and
$\Phi_{b_{i}+\delta_{i}\sigma^{2}/c,\sigma^{2},A_{i}}(Y)=\alpha/2.$ (3.6)
The confidence interval
$\mathrm{CI}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}(Y):=\left[a_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)},b_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right]$ (3.7)
has $1-\alpha$ conditional coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(Y+\zeta,\delta\right).$
Computing (3.7) requires numerically solving (3.5) for $a_{i}$ and (3.6) for $b_{i}$, where $i=\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$.
Thus, both (3.4) and (3.7) provide confidence intervals for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that attain $1-\alpha$ unconditional coverage. We should prefer the approach that yields narrower intervals. Figure 1(a) displays the ratio of the widths of the two intervals when $\mu=0$ and $\sigma^{2}=1$, as we vary the sample size $n$ and the variance $2c^{2}$ of the Laplace noise in (3.3).
The infer-and-widen interval (3.4) involves tuning parameters $\nu$ and $\eta$; to construct Figure 1(a), for a fixed value of the noise $c$ and sample size $n$ in (3.3), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. The data fission intervals (3.7) are much narrower than the infer-and-widen intervals (3.4), across all values of $n$ and $c$ considered.
It is possible that the width of the infer-and-widen interval (3.4) is due to the fact that [28] apply a series of inequalities in its derivation; furthermore, it is natural to wonder whether another infer-and-widen proposal could perform better. Therefore, we now consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $Y_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that achieves a given coverage. It is not computable in practice, as it requires knowledge of the true parameter value $\mu$. In Figures 1(b, c), we numerically compare the “oracle” infer-and-widen method to the data fission interval (3.7) under the selection rule (3.3) with variances $2c^{2}=0.33$ and $3$, respectively. The former method has a wider average width than the latter: thus, no infer-and-widen interval can outperform the data fission interval (3.4) in this setting.
Section C of the supplementary materials contains a related comparison of methods for inference after non-randomized selection using 3.2. The conclusions are the same.
Figure 1: Vignette #1 under (3.1) with $\mu=0$, $\sigma=1$, and the selection rule in $3.$ Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of the infer-and-widen (I&W) 3.4 and data fission (DF) 3.7 intervals with $95\%$ coverage across values of the sample size $n$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. (b, c): Fixing $n=100$, we display the average width of the “oracle” infer-and-widen interval and the data fission interval 3.7 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval based on algorithmic stability 3.4 are also displayed. The Laplacian noise variance is low and high in panels (b) and (c), respectively.
4 Vignette #2: Maximal contrasts
Following Zrnic and Jordan [28], we consider the model 3.1 and suppose we also have access to a fixed design matrix with normalized columns $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Our interest lies in the quantity $X_{\mathcal{S}(Y)}^{\top}\mu$ defined via the selection rule
$\mathcal{S}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X_{j}^{\top}Y\right|;$ (4.1)
[28] refer to this as the “maximal contrast”. If $X=I$, then this setting is similar to Vignette #1, except with an absolute value in the selection rule.
To achieve a “stable” algorithm, [28] consider the selection rule
$\mathcal{S}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X^{\top}_{j}Y+\zeta_{j}\right|,$ (4.2)
where $\zeta_{1},\ldots,\zeta_{p}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$ and $c$ is a positive constant that governs the scale of the additive noise.
Proposition 3 (An infer-and-widen interval based on (4.2), given by Zrnic and Jordan [28]).
Under (3.1), consider the Laplacian selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ defined in (4.2). For any $\eta>0$ and $\nu\in(0,1)$ such that $c\geq 2z_{1-\alpha\nu/(2p)}/\eta$, the infer-and-widen confidence interval
$\mathrm{CI}_{{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\nu)e^{-\eta}}(Y):=\left[X^{\top}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}Y\pm\sigma z_{1-\alpha(1-\nu)e^{-\eta}/2}\right]$ (4.3)
has $1-\alpha$ unconditional coverage for $X^{\top}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\mu$ in the sense of 2.1.
We will take a randomized conditional inference approach [24, 20] to derive an alternative confidence interval that makes use of the identical selection event.
Proposition 4 (A randomized conditional selective inference interval based on 4.2).
Consider the model in 3.1 and selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ in 4.2. Define
$W$ $:=(I_{n}-X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top})Y,$
$\delta$ $:={\rm sign}\left(X^{\top}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}Y+\zeta_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right),$
and $a_{j}$ and $b_{j}$ to satisfy
$\Phi_{a_{j},\sigma^{2},[v_{\min},v_{\max}]}(X_{j}^{\top}Y)=1-\alpha/2$ (4.4)
and
$\Phi_{b_{j},\sigma^{2},[v_{\min},v_{\max}]}(X_{j}^{\top}Y)=\alpha/2,$ (4.5)
where $\Phi_{\lambda,\sigma^{2},[v_{\min},v_{\max}]}(\cdot)$ is defined in Proposition 2, and $v_{\min}$ and $v_{\max}$ are defined in Appendix A. The confidence interval
$\mathrm{CI}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}\left(Y\right):=\left[a_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)},b_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right]$ (4.6)
has $1-\alpha$ conditional coverage for $X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(\mathcal{S}_{c}^{\mathrm{Laplace}}(Y),\zeta,W,\delta\right)$.
Computing 4.6 requires numerically solving 4.4 for $a_{i}$ and 4.5 for $b_{i}$, where $i=\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$.
We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X\phi$ where half of the elements of $\phi$ are zero and the other half are drawn from an exponential distribution with mean $50/7$. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and its columns are subsequently scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Figure 2(a) displays the ratio of confidence interval widths across values of $p$ and the variance of the Laplacian noise $2c^{2}$ in 4.2. The infer-and-widen interval (4.6) involves tuning parameters $\nu$ and $\eta$; to construct Figure 2(a), for a fixed value of the noise $c$ and number of features $p$ in (4.2), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 3. The randomized conditional selective inference intervals are narrower than the infer-and-widen intervals across most values of $p$ and $c$ considered.
Once again, we consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top}Y$ that achieves a given coverage. In Figures 2(b, c), we numerically compare the “oracle” infer-and-widen method to the randomized conditional selective inference interval (4.6) under the selection rule (4.2) with variances $2c^{2}=0.33$ and $3$, respectively. In this setting, the “oracle” infer-and-widen interval is narrower than the randomized conditional selective inference interval in (4.6), thereby indicating that in principle, an infer-and-widen interval can outperform the randomized conditional selective inference interval (4.6) in this setting. However, to the best of our knowledge, no methods to construct such a narrower infer-and-widen interval are currently available.
Figure 2: Vignette #2 under (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in 4.2, where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of infer-and-widen (I&W) 4.3 and randomized conditional selective inference (RCSI) 4.6 intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. (b, c): Fixing $p=100$, we display the average width of the “oracle” infer-and-widen interval and the RCSI interval 4.6 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval 4.3 based on algorithmic stability are also displayed. The Laplacian noise variance is low and high in panels (b) and (c), respectively.
5 Vignette #3: Inference after the lasso
We now consider inference on the set of coefficients selected by the lasso in a linear model. We again consider the model 3.1, alongside a fixed design matrix $X\in\mathbb{R}^{n\times p}$. Our selection rule is the set of features selected by the lasso estimator $\hat{\beta}^{(\lambda)}$ [25]:
$\mathcal{S}(Y):={\bf{\rm supp}}\;\left(\hat{\beta}^{(\lambda)}\right),$ (5.1)
where
$\hat{\beta}^{(\lambda)}:=\mathop{\mathrm{arg\,min}}_{\beta}\left\{\tfrac{1}{n}\left\|Y-X\beta\right\|^{2}_{2}+\lambda\|\beta\|_{1}\right\}.$
This is a “non-randomized” selection event, and we defer to Section E of the supplement for inference after a related randomized selection event.
Our goal is valid inference on the least squares coefficients in the model selected by $\mathcal{S}(Y)$:
$\beta_{\mathcal{S}(Y)}:=\left(X_{\mathcal{S}(Y)}^{\top}X_{\mathcal{S}(Y)}\right)^{-1}X_{\mathcal{S}(Y)}^{\top}\mathbb{E}\left[Y\right],$
where $X_{\mathcal{S}(Y)}$ is the subset of columns of $X$ indexed by $\mathcal{S}(Y)$. Since $\mathcal{S}(Y)$ is a data-adaptive model, the naive point estimator
$\hat{\beta}_{\mathcal{S}(Y)}:=\left(X_{\mathcal{S}(Y)}^{\top}X_{\mathcal{S}(Y)}\right)^{-1}X_{\mathcal{S}(Y)}^{\top}Y$
is no longer guaranteed to be unbiased for $\beta_{\mathcal{S}(Y)}$.
Obtaining valid inference on $\beta_{\mathcal{S}(Y)}$ in the sense of 1.2 is a classic problem in the field of selective inference, and there are two main categories of methods which do so. The first category guarantees unconditional coverage, 1.2, via the stronger goal of simultaneous coverage. Since the Bonferroni correction [8] is valid but would require an overly conservative adjustment for all $O(2^{p})$ possible models, Berk et al. [3] derive the narrower PoSI simultaneous correction. Zrnic and Fithian [26] subsequently developed the locally simultaneous inference approach which focuses the simultaneous correction on only a subset of likely models, thereby often improving power; these two approaches yield infer-and-widen intervals centered at $\hat{\beta}_{\mathcal{S}(Y)}$ and with widths proportional to the diagonal elements of $\left(X_{\mathcal{S}(Y)}^{\top}X_{\mathcal{S}(Y)}\right)^{-1/2}$ with a shared constant of proportionality. The second category of methods approach unconditional coverage via the stronger goal of conditional coverage, 1.3. The foundational work of Lee et al. [16] characterized the selection event $\mathcal{S}(Y)$, and used this to derive the conditional distribution $\hat{\beta}_{\mathcal{S}(Y)}\mid\mathcal{C}(Y)$ for inference on $\beta_{\mathcal{S}(Y)}$, where the sigma algebra of $\mathcal{C}(Y)$ is a superset of that of ${\mathcal{S}}(Y)$. In order to improve upon the power of this method, McCloskey [18] developed the hybrid method, which uses the conditional confidence interval in cases where there is strong signal, and otherwise resorts to unconditional intervals; thus the hybrid method provides only unconditional and not conditional coverage.
As before, we compare the widths of these methods’ confidence intervals, as well as the “oracle” infer-and-widen interval: the narrowest possible interval centered at $\hat{\beta}_{\mathcal{S}(Y)}$ and with width proportional to the diagonal elements of $\left(X_{\mathcal{S}(Y)}^{\top}X_{\mathcal{S}(Y)}\right)^{-1/2}$ with a shared constant of proportionality that attains coverage at a given level. We simulate data according to 3.1 with $n=100$, $p=10$, $\sigma=1$, and fixed penalty $\lambda$. We set $\mu=X\beta$, where $\beta_{1}=\dots=\beta_{5}$ and $\beta_{6}=\dots=\beta_{10}=0$. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with correlation $\rho$. Figure 3(a) displays the ratio of average $95\%$ confidence interval widths of the locally simultaneous inference (LSI) [26] and hybrid [18] methods across values of $\rho$ and $\beta_{1}/\lambda$. A ratio greater than one indicates that the LSI method has a wider average interval width. In Figure 3(b, c), we fix $\rho=0.9$, and set $\beta_{1}/\lambda=0$ in panel (b) and $\beta_{1}/\lambda=1$ in panel (b). We numerically compare the “oracle” interval to the conditional [16] and hybrid [18] intervals. For reference, we also display the average empirical widths and coverages of the $95\%$ intervals from the classic, simultaneous PoSI (SI) [3], and LSI [26] approaches. In this setting, the “oracle” infer-and-widen interval is the narrowest interval, thereby indicating that in principle, an infer-and-widen interval can outperform the hybrid interval. However, the hybrid method outperforms both existing infer-and-widen methods.
Figure 3: Vignette #3 under (3.1) with $n=100$, $p=10$, $\sigma=1$, $\mu=X\beta$, and the selection rule in 5.1, where the rows of $X$ are equi-correlated normally distributed, and $\beta$ has five nonzero elements. Results are averaged over $100$ simulated datasets. (a): We display the ratio of the average widths of $95\%$ confidence intervals from the LSI infer-and-widen method [26] and the hybrid method [18]. A width ratio that exceeds one implies that the infer-and-widen interval is wider. (b, c): Fixing $\rho=0.9$, we compare the average width of the “oracle” infer-and-widen (I&W) interval to the conditional [16] and hybrid [18] intervals. For reference, we also display the average empirical widths and coverages of the $95\%$ intervals from the classic, simultaneous PoSI (SI) [3], and LSI [26] approaches. Panels (b) and (c) contain zero and nonzero signal, respectively.
6 The bias of infer-and-widen intervals
We now investigate why infer-and-widen intervals are too wide. To facilitate the analysis, we consider Vignette #1 with a simple selection rule that uses Gaussian noise,
$\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}(Y_{i}+c\zeta_{i}),\;\;\;\;\zeta_{i}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm N}}\,(0,\sigma^{2}).$ (6.1)
We know from Section 2 that any infer-and-widen interval is centered around $Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}$.
Proposition 5 (Bounding the bias of the infer-and-widen midpoint).
Under (3.1), with $\mu_{1}=\dots=\mu_{n}$, consider the selection rule $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)$ in (6.1). Then
$\mathbb{E}\left[Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]\geq\frac{\sigma}{2}(\sqrt{1+c^{2}}-c)\sqrt{\log n}.$
That is, the bias increases on the order of $\sqrt{\log n}$. Note that as $c\to\infty$, i.e., as the selection rule becomes fully random, this lower bound goes to zero as expected.
Noting that $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)$ in (6.1) is independent of $Y-\frac{1}{c}\zeta$ [19], it is straightforward to show that $(Y-\frac{1}{c}\zeta)_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\pm\sigma(1+c^{-2})^{1/2}z_{1-\alpha/2}]$ is a valid $1-\alpha$ conditional confidence interval for $\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}$ in the sense of 1.3. Furthermore,
$\mathbb{E}\left[\left(Y-\frac{1}{c}\zeta\right)_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]=0,$
i.e., the midpoint is unbiased for the mean of the selected observation.
Thus, infer-and-widen here leads to a confidence interval that is centered in the wrong place, but a simple alternative that makes use of the same selection event yields an interval that is correctly centered.
7 Discussion
In this paper, we compared the performance of the “infer-and-widen” approach to simple alternative methods in three vignettes: two taken from [28], and the third involving feature selection via the lasso. Our findings are as follows:
  1. 1.
    In Vignette #1, the “oracle” interval was wider than the alternative approaches we consider; thus, there may be limited value in developing narrower infer-and-widen methods that can be computed in practice.
  2. 2.
    In Vignettes #2 and #3, the “oracle” interval is the narrowest interval, but infer-and-widen intervals that can be computed in practice are often much wider.
One may wonder whether our findings conflict with those of Goeman and Solari [13], who proved that any method that guarantees conditional coverage, (1.3), is dominated by an unconditional multiple testing procedure defined on the full universe of hypotheses. However, their setting is quite different from ours: they consider inference on non-selected parameters and unconditional procedures outside of the infer-and-widen framework.
Appendix A Deferred notation
In the context of Proposition 4 , we define
$A:=\begin{bmatrix}0_{1\times n_{1}}&-1&0_{1\times n_{2}}\\ I_{n_{1}}&-1_{n_{1}}&0_{n_{1}\times n_{2}}\\ 0_{n_{2}\times n_{1}}&-1_{n_{2}}&I_{n_{2}}\\ -I_{n_{1}}&-1_{n_{1}}&0_{n_{1}\times n_{2}}\\ 0_{n_{2}\times n_{1}}&-1_{n_{2}}&-I_{n_{2}}\end{bmatrix}$,
where we denote $n_{1}:=\mathcal{S}_{c}^{\mathrm{Laplace}}(y;\zeta)-1$ and $n_{2}:=p-\mathcal{S}_{c}^{\mathrm{Laplace}}(y;\zeta)$ for brevity. Then we can define
$v_{j}$ $:=-(A\zeta+AX^{\top}W)_{j}/(AX^{\top}X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)})_{j},$
$v_{\max}$ $:=\underset{j:\left(\delta AX^{\top}X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right)_{j}<0}{\mathrm{max}}v_{j},$
$v_{\min}$ $:=\min_{j:\left(\delta AX^{\top}X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right)_{j}>0}v_{j}.$
{funding}
R.P. was partially supported by an Amazon AI research award. D.W. acknowledges funding from NSF DMS 2322920, NSF DMS 2514344, NIH 5P30DA048736, ONR N00014-23-1-2589, and a Simons Investigator Award for Mathematical Modeling of Living Systems.
[…]